OpenStudy (itsmichelle29):

What is the quotient of 1 + i and its conjugate? i -1 -2i 2

OpenStudy (alexandervonhumboldt2):

for a+bi a-bi is a conjugate

OpenStudy (alexandervonhumboldt2):

so we have (1+i)*(1-i)=1*1-1*i+1*i-(i*i)=1-(i*i)=??? remember: i*i=-1

OpenStudy (itsmichelle29):

2???? Yes

OpenStudy (alexandervonhumboldt2):

correct

OpenStudy (victoriasushchik):

yes thats correct!

OpenStudy (itsmichelle29):

thanks